x is normally distributed with a mean of 100 and a standard deviation of 20. what is the probability that x is greater than 145
1
Expert's answer
2019-09-20T10:27:12-0400

The probability density function for normal (gaussian) distribution is


"\frac{1}{{\sigma \sqrt {2\pi } }}{e^{ - \frac{{{{(x - \mu )}^2}}}{{2{\sigma ^2}}}}}"

In this case mean "\mu = 100" and standard deviation "\sigma = 20" .

1)Use standard normal variable Z


"Z = \frac{{X - \mu }}{\sigma } \Rightarrow X = \sigma Z + \mu"

so the probability that the random variable X is greater than 145 can be represented in standard form


"P(X > 145) = P(20Z + 100 > 145) = P(Z > 2.25)"


Now we can use standard normal distribution table (Z-table) (I get it from z-table.com)


It represents the probability that "Z < a", so we need to use the normalized property (full probability is equal to 1), so "P(Z > a) = 1 - P(Z < a)". In our case from table "P(Z < 2.25) \approx 0.9878" so



"P(X > 145) = P(Z > 2.25) = 1 - P(Z < 2.25) = 1 - 0.9878 = 0.0122"

The answer is 0.0122.

Answer: "P(X > 145) \approx 0.0122"